# Narrow linewidth exciton-polariton laser

### v1

@@ 5 Fluctuations and time scale analysis
Finally, we analyse the effect of fluctuations in energy observed here and its relation to the measured linewidth. At long timescales ($>10^{-4}$ s), longer than the time required to scan over a single spectral peak, fluctuations can lead to a "slow drift" of the condensate energy. The scanning of the SFPI over 3 FSRs allows us to measure the energy drift after 2.5 ms. An example distribution of the energy shifts is shown in Fig. 2(h), showing a mean of $E_{shift}/h\approx 1.0$ GHz. The former can arise from a slow but consistently increasing excitation power within the 8-ms quasi-CW pulse, which is due to increased AOM efficiency resulting from transient heating. Heating of the sample due to the excitation laser can also contribute but we expect a redshift instead of blueshift observed here.
The fluctuations in energy, quantified by the standard deviation $\Delta E$, are $\Delta E=0.64$ GHz within the shot (every 2.5 ms) and $\Delta E=0.80$ GHz from shot to shot (every 800 ms), for the dataset in Fig. 2(h). The former is generally smaller than the latter. For example, the $\Delta\ell_{\nu}$ for each power in Fig. 2(c) averages around $\Delta E\approx 1.0$ GHz and $\Delta E\approx 1.6$ GHz, respectively. This means that there is less energy drift (or less fluctuations) on the ms time scale compared to longer time scale. Furthermore, the energy fluctuations do not have a strong dependence on the excitation power, trap offset, and ring diameter (see SI for the relevant plots).
We can estimate the excitation power fluctuations responsible for the observed energy fluctuations using the slope of the linear fits on the experimental $E(P)$ in Fig. 2(b). Fluctuations around $\Delta E=1$  GHz correspond to power fluctuations of 22% and 4% for $P/P_{th}<1.3$ and $P/P_{th}>1.3$ in Fig. 2(b), respectively. The latter is comparable to the measured intensity fluctuations of the laser in this timescale. The former is quite large and can arise from other sources such as SLM flicker, air currents, and mechanical vibrations of the cryostat, which can all lead to large intensity modulation and variations in exciting the same spot on the sample. In principle, an active feedback could be implemented using the AOM to modulate the excitation intensity to minimize the energy drift at these timescales.
@@
To probe the faster timescales ($<10^{-4}$ s), we intentionally modulate the excitation laser intensity using the AOM to externally introduce controllable fluctuations. Sample spectra are presented in Fig. 4(a) for different amplitudes of a 1-MHz sinusoidal modulation. The results for low frequencies are presented in the SI. The lineshape of the spectra are clearly affected by the added modulation such that it deviates from a Lorentzian to a flat-top and eventually towards a two-peak profile. We define the effective linewidth using the FWHM of the spectra and their dependence on the modulation strength is presented in Fig. 4(b), clearly showing the broadening of the spectra with increasing modulation strength. This suggests that if fast modulations in $P$ or in the reservoir density are present, they naturally lead to broadening of the linewidth.
@@
Finally, we demonstrate the extreme sensitivity of the linewidth on small shot-to-shot fluctuations in energy $E$ or power $P$ close to the threshold. The simultaneous energy and linewidth measurements taken from single quasi-CW pulses of the condensate are plotted in Fig. 4(e) for the experimental conditions corresponding to the 6-$\mu$m trap in Fig. 3(d), which has an average $P/P_{th}\approx 1.1$. The plot clearly shows that if the energy blueshifts by a small amount $\sim$1 GHz, the linewidth narrows by an order of magnitude. This means that the large measured fluctuations in linewidth $\Delta\ell_{\nu}$ close to the threshold are due to this extreme sensitivity. At higher $P/P_{th}$, when the linewidth plateaus, the $E$-$\ell_{\nu}$ correlation vanishes [see for example Fig. 2(i)] the measured $\Delta\ell_{\nu}$ reflects the uncorrelated fluctuations in linewidth.
@@ 6 Discussions
We presented precise linewidth and energy shift measurements of the laser emission from an optically trapped single-mode polariton condensate using a scanning Fabry-Pérot interferometer. We found three general behaviors regarding the linewidth. First, the linewidth drastically narrows with increasing excitation power and plateaus to a median of $\ell_{\nu}<100$ MHz for $P/P_{th}>1.2$. This behavior is independent of the overlap with the reservoir, as confirmed by increasing either the trap offset or the trap size, as long as the condensate is trapped. Second, the reservoir only strongly affects the linewidth at excitation powers close to threshold ($P/P_{th}<1.2$), where larger overlap leads to larger linewidths. Third, the fluctuations in the measured linewidth decrease with excitation power down to $\Delta\ell_{v}=8$ Mhz at $P/P_{th}\approx 1.8$. Hence, to optimize the linewidth, the normalized excitation power should be increased beyond $P/P_{th}>1.2$, followed by reduction of the overlap of the condensate with the reservoir.
To compare our results with previous works, we converted the reported coherence times to linewidths using Eqs. (1) and (2). The narrowest linewidth we measured ($\ell_{\nu}=56$ MHz) is four times narrower than the previous record [18], but note that the latter was only extrapolated from a limited dataset. In terms of coherence time, the Lorentzian lineshape in our case translates to $\tau_{coh}=1/{\pi\ell_{\nu}}=5.7$ ns, which is twice longer than in Ref. [18]. The comparison with other reported values [25, 16, 17, 24, 22, 18] of coherence time and linewidths are tabulated in the SI.
@@
Our work provides valuable experimental insights into the spectral properties of the polariton laser emission and the condensate itself. We clearly observe energy fluctuations $\Delta E$ on the order of 1 Ghz on timescales greater than 10-3 s, independent of the experimental conditions. This drift can be minimised or controlled via active feedback using the AOM, similar to atomic condensates [2]. Faster feedback towards the GHz regime using an electro-optic modulator may result in an even smaller linewidth.

### v2

@@ 5 Fluctuations and time scale analysis
Finally, we analyze the fluctuation in the condensate energy $E$ and its relation to the measured linewidth $\ell_{\nu}$. Fluctuation in energy leads to linewidth broadening, but the energy fluctuation and linewidth measurements occur at different timescales. It takes about 2.5 ms to scan over a single FSR or to measure any energy shift, but it only takes $<50\leavevmode\nobreak\ \mu$s to scan over a single narrow spectral peak. Hence, we can probe effects of slow and fast noise by probing the fluctuation in $E$ and the linewidth, respectively.
We first look at the slow fluctuation by analyzing the distribution of the energy shifts, such as the one shown in Fig. 2(h), exhibiting a mean shift of $\approx 1.0$ GHz. The mean shift can arise from a slow increase in excitation power within the 8-ms quasi-CW pulse due to a slow increase in AOM efficiency after turning on. Heating of the sample due to the excitation laser can also contribute, but would result in a redshift instead of the blueshift observed here.
The fluctuation above the mean energy shift, quantified by the standard deviation $\Delta E$, is $\Delta E=0.64$ GHz within the shot (every 2.5 ms) and $\Delta E=0.80$ GHz from shot to shot (every 800 ms), for the dataset in Fig. 2(h). The former is generally smaller than the latter such that for each power in Fig. 2(c) the average is around $\Delta E\approx 1.0$ GHz and $\Delta E\approx 1.6$ GHz, respectively. This means that there is less energy drift (or less fluctuation) on the ms time scale compared to longer time scale. Furthermore, the energy fluctuation does not have a strong dependence on the excitation power, trap offset, and ring diameter (see Supplementary Document for the relevant plots).
Assuming the excitation power is the only source of noise, we can estimate the equivalent fluctuations in $P$ responsible for the observed energy fluctuation using the slope of the linear fits on the experimental $E(P)$ in Fig. 2(b). Fluctuation of $\Delta E\approx 1$  GHz correspond to power fluctuations of 22% and 4% for $P/P_{th}<1.3$ and $P/P_{th}>1.3$ in Fig. 2(b), respectively. The latter is comparable to the measured intensity fluctuations of the laser in this timescale. The former is quite large and can arise from the other sources of technical noise such as SLM flicker, air currents, and mechanical vibrations of the cryostat, which can all lead to large intensity modulation and shifts the excitation spot on the sample. The latter can change the excitation profile and the polariton detuning resulting in changing $P_{th}$ and polariton energy $E_{0}$ (see Eq. 3), cumulatively changing the condensate energy $E$ even if the input power $P$ is the same. An active feedback can be implemented using the AOM to modulate the excitation intensity to minimize the energy drift at these timescales.
@@
To probe the faster timescales ($<10^{-4}$ s), we analyze the linewidth by intentionally modulating the excitation laser intensity using the AOM to externally introduce controllable fluctuations. Sample spectra are presented in Fig. 4(a) for different amplitudes of a 1-MHz sinusoidal modulation. The results for low frequencies are presented in the SI. The lineshape of the spectra are clearly affected by the added modulation such that it deviates from a Lorentzian to a flat-top and eventually towards a two-peak profile. We define the effective linewidth using the FWHM of the spectra and their dependence on the modulation strength is presented in Fig. 4(b), clearly showing the broadening of the spectra with increasing modulation strength. This suggests that if fast modulations in $P$ or in the reservoir density are present, they naturally lead to broadening of the linewidth.
@@
Finally, we demonstrate the extreme sensitivity of the linewidth on small shot-to-shot fluctuations in energy $E$ or power $P$ close to the threshold. The simultaneous energy and linewidth measurements taken from single quasi-CW pulses of the condensate are plotted in Fig. 4(e) for the experimental conditions corresponding to the 6-$\mu$m trap in Fig. 3(d), which has an average $P/P_{th}\approx 1.1$. The plot clearly shows that if the energy blueshifts by a small amount $\sim$1 GHz, the linewidth narrows by an order of magnitude. This means that the large measured fluctuations in linewidth $\Delta\ell_{\nu}$ close to the threshold are due to this extreme sensitivity. At higher $P/P_{th}$, when the linewidth plateaus, the $E$-$\ell_{\nu}$ correlation vanishes [see for example Fig. 2(i)] and the measured $\Delta\ell_{\nu}$ reflects the uncorrelated fluctuations in linewidth.
@@ 6 Discussions
We presented precise linewidth and energy shift measurements of the laser emission from an optically trapped single-mode polariton condensate using a scanning Fabry-Pérot interferometer. We found three general behaviors regarding the linewidth. First, the linewidth drastically narrows with increasing excitation power and plateaus to a median of $\ell_{\nu}<100$ MHz (measured within 50 $\mu$s) for $P/P_{th}>1.2$. This behavior is independent of the overlap with the reservoir, as confirmed by increasing either the trap offset or the trap size, as long as the condensate is trapped. Second, the reservoir only strongly affects the linewidth at excitation powers close to threshold ($P/P_{th}<1.2$), where larger overlap leads to larger linewidths. Third, the fluctuations in the measured linewidth decrease with excitation power down to $\Delta\ell_{v}=8$ Mhz at $P/P_{th}\approx 1.8$. Hence, to optimize the linewidth, the normalized excitation power should be increased beyond $P/P_{th}>1.2$, followed by reduction of the overlap of the condensate with the reservoir.
To compare our results with previous works, we converted the reported coherence times to linewidths using Eqs. (1) and (2). The narrowest linewidth we measured ($\ell_{\nu}=56$ MHz) is four times narrower than the previous record [18], but note that the latter was only extrapolated from a limited dataset. In terms of coherence time, the Lorentzian lineshape in our case translates to $\tau_{coh}=1/{\pi\ell_{\nu}}=5.7$ ns, which is twice longer than in Ref. [18]. The comparison with other reported values [25, 16, 17, 24, 22, 18] of coherence time and linewidths are tabulated in the Supplemental Document.
@@
Our work provides valuable experimental insights into the spectral properties of the polariton laser emission and the condensate itself. We clearly observe energy fluctuations $\Delta E$ on the order of 1 GHz on timescales greater than 10-3 s, independent of the experimental conditions. This drift can be minimised or controlled via active feedback using the AOM, similar to atomic condensates [2]. Faster feedback towards the GHz regime using an electro-optic modulator may result in an even smaller linewidth. Electrical injection [46, 47] of polariton lasers can provide a suppression of the large $\Delta E$ at low frequencies since the pump and the cavity are embedded in one chip. However, it may also introduce a new set of noise. Further research in this area would be valuable.
@@
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]